OpenStudy (anonymous):

7. Micronet dropped to $3.60 a share in October. What was the percent of decrease from July's per share price? 8. By the end of the year, Micronet's price rose back to $4.84 per share. What is the percent of increase from its original price?

OpenStudy (turingtest):

You posted twice apparently... 5) 4.4+4.4x=6.82 4.4(1+x)=6.82 x=6.82/4.4-1=0.55=55% 6) 6.82-6.82x=5.56 6.82(1-x)=5.56 x=1-5.56/6.82=0.18=18% I'll set up the last two for you, you can solve them yourself: 7) 5.56-5.56x=3.6 8) 4.4+4.4x=4.84

OpenStudy (turingtest):

or can you really not solve these last two for x?